### v1

# Déjà Vu: Efficient Video-Language Query Engine with Learning-based Inter-Frame Computation Reuse

Abstract.
Recently, Video-Language Models (VideoLMs) have demonstrated remarkable capabilities, offering significant potential for flexible and powerful video query systems. These models typically rely on Vision Transformers (ViTs), which process video frames individually to extract visual embeddings. However, generating embeddings for large-scale videos requires ViT inferencing across numerous frames, posing a major hurdle to real-world deployment and necessitating solutions for integration into scalable video data management systems. This paper introduces Déjà Vu, a video-language query engine that accelerates ViT-based VideoLMs by reusing computations across consecutive frames. At its core is ReuseViT, a modified ViT model specifically designed for VideoLM tasks, which learns to detect inter-frame reuse opportunities, striking an effective balance between accuracy and reuse. Although ReuseViT significantly reduces computation, these savings do not directly translate into performance gains on GPUs. To overcome this, Déjà Vu integrates memory-compute joint compaction techniques that convert the FLOP savings into tangible performance gains. Evaluations on three VideoLM tasks show that Déjà Vu accelerates embedding generation by up to a 2.64$\times$ within a 2% error bound, dramatically enhancing the practicality of VideoLMs for large-scale video analytics.
PVLDB Reference Format:
PVLDB, 18(1): XXX-XXX, 2025.
doi:XX.XX/XXX.XX This work is licensed under the Creative Commons BY-NC-ND 4.0 International License. Visit https://creativecommons.org/licenses/by-nc-nd/4.0/ to view a copy of this license. For any use beyond those covered by this license, obtain permission by emailing info@vldb.org. Copyright is held by the owner/author(s). Publication rights licensed to the VLDB Endowment.
Proceedings of the VLDB Endowment, Vol. 18, No. 1 ISSN 2150-8097.
doi:XX.XX/XXX.XX
PVLDB Artifact Availability:
The source code, data, and/or other artifacts have been made available at https://github.com/casys-kaist/DejaVu.
1. Introduction
Video data is pervasive in today’s world, playing a critical role in various applications such as autonomous machines (Deb et al., 2018; Gawande et al., 2020; Hadidi et al., 2018), education (Gan et al., 2023), healthcare (Sanderson and Matuszewski, 2022), security and surveillance (Li et al., 2019; Toor et al., 2019), e-commerce (Zhang et al., 2020a, 2022), and many others (Mun et al., 2017; Li et al., 2022d). Recently, Video-Language Models (VideoLMs) have demonstrated remarkable capabilities, offering significant potential for flexible and powerful video analytics tasks (Colas et al., 2019; Gupta and Demner-Fushman, 2022; Zhao et al., 2023; Ju et al., 2022; Yang et al., 2021, 2022a). While these powerful algorithmic advances promise enormous potential, the massive computational demand poses a significant hurdle for the widespread adoption of VideoLMs for video query systems.
The excessive computational demand arises from two primary factors. First, modern VideoLMs heavily rely on Vision Transformers (ViTs) (Dosovitskiy et al., 2021) as their core engine for visual feature extraction. ViTs contain hundreds of millions of parameters and require tens of billions of FLOPs for each inference. Second, ViT inference must be performed iteratively across numerous video frames, which can be exceptionally large in number (e.g., a 1-hour video sampled at 2 FPS contains 7,200 frames). Consequently, utilizing VideoLMs for large-scale video analytics becomes nearly infeasible, even with multi-GPU datacenter resources.
Figure 1. Overview of video language query systems supporting three example VideoLMs. Modern VideoLMs employ vision transformer (ViT) as their core engine.
Figure 1 provides an overview of video-language query systems, where ViTs process video frames to extract visual embeddings, which are then fed into query-specific operations such as video retrieval, video question answering, and video question grounding. The critical role of ViTs in this process emphasizes the importance of improving ViT inference efficiency to unleash the potential of VideoLMs on building video-language query systems.
Predictably, the vital need for efficient ViT inferencing has led to a large body of acceleration work in recent years (Rao et al., 2021; Fayyaz et al., 2022; Liang et al., 2022; Kong et al., 2022; Chen et al., 2023; Cao et al., 2023; Li et al., 2022a; Wang et al., 2021b; Dong et al., 2023; Qin et al., 2023; Bolya et al., 2023; Marin et al., 2021). Most of these works focus on intra-frame computation reuse, targeting inter-token feature similarities since they aim to speed up a single ViT inference run processing one image input. While effective for single images, these methods overlook additional opportunities present in video data that exhibit inter-frame similarities across multiple consecutive frames.
Recently, several works (Song et al., 2024; Dutson et al., 2023; Choi et al., 2024) have pioneered the exploitation of inter-frame computation reuse for ViT inferencing on video applications. However, these works have the following two limitations, which leave them in a position of questionable utility for most practical use cases:
  • Limitation 1: These methods manually identify reuse opportunities for given models and fixedly enable computation reuse, making it considerably difficult to locate the sweet spot on the accuracy-reuse tradeoff space.
  • Limitation 2: Although these techniques enable significant FLOP reduction, they do not yield corresponding performance gains, as the remaining computations are sparse and, therefore, inefficient on GPUs.
To address these limitations, this work proposes Déjà Vu, an algorithm-system co-designed query engine that leverages a learning-based approach to automatically identify the computation reuse opportunities and find the sweet spot in the tradeoff space between accuracy and reuse.
The key challenges are (1) to maximize the reuse opportunities without causing significant accuracy loss in VideoLM applications and (2) to effectively translate the FLOPs savings into performance gains. Déjà Vu tackles these challenges by making the following two key contributions, each corresponding to the two limitations:
(1):
ReuseViT: Reuse-enabling ViT model. We develop a modified ViT model, referred to as ReuseViT, that reuses precomputed values for similar tokens between frames. To maximize reuse opportunities, we devise a frame referencing strategy that determines inter-frame reference chains for computation reuse. ReuseViT contains Decision Layers that dynamically determine when to reuse computations by processing multiple inputs such as cosine similarity, attention scores, codec metadata, and reference types, allowing it to learn the importance and interactions of these hints automatically. Additionally, it employs Restoration Layers that learn to calibrate the reused computation values based on the changes occurred in the current frame.
To train ReuseViT, we use reparameterization with soft gating via the Gumbel-Softmax mechanism, enabling backpropagation through discrete decisions. This allows the model to mimic hard gating, allowing gradients to flow. Further, we model error propagation by training with grouped frames to mitigate error accumulation from reusing computations across multiple frames.
(2):
Memory and compute compactions for fast and efficient ReuseViT inferencing. To achieve significant performance gains from these algorithmic innovations, we propose three key system-level techniques. (1) We introduce layer-wise computation scheduling, which processes computations for multiple frames in a layer-by-layer manner. (2) This approach enables cached memory compaction, a memory management technique that clears caches after each layer is completed, allowing for optimized memory usage. Consequently, this increases the batch size during inference, leading to better hardware utilization. (3) Additionally, we implement sparse computation compaction, which restructures irregular data patterns caused by computation reuse into dense computations suitable for efficient GPU execution. By consolidating tokens from multiple frames, we create more regular matrices, improving the efficiency of matrix multiplication operations on GPUs, particularly when reuse rates are high.
To demonstrate the effectiveness of Déjà Vu, we perform evaluations using three different VideoLMs – (1) Video Retrieval, (2) Video Question Answering, and (3) Video Question Grounding – on their respective datasets. We observe that Déjà Vu achieves throughput improvement of 1.81$\times$, 2.64$\times$, and 2.54$\times$ for the three tasks, respectively, within 2% error bound. Under the same accuracy constraints, the state-of-the-art systems using inter-frame computation reuse offer only 1.32$\times$, 2.08$\times$, 2.20$\times$ speedups for the respective tasks.
These significant performance gains, coupled with minimal accuracy loss, underscore the solution’s potential to bridge the computational gap, paving the way for VideoLMs to unlock new capabilities.
2. Background and Motivation
2.1. Video Query Processing
A rich body of research explores accelerating large-scale video queries using AI models. We categorize them into three groups.
Task-specific CNN pipelines. Early approaches (Kang et al., 2017; Kang et al., VLDB; Lai et al., 2022; Cao et al., 2022) accelerate queries by training small, query-specific models as approximations for computationally expensive deep learning pipelines. Systems such as NoScope (Kang et al., 2017) and BlazeIt (Kang et al., VLDB) train lightweight classifiers or regression models tailored to each query, significantly reducing inference costs for known object classes. However, these methods are inherently inflexible, as they require model training or selection for every new query, making them unsuitable for quickly adapting to new or changing requirements.
Task-agnostic proxy embeddings. The second category (Kang et al., 2021; Bang et al., 2023) precomputes a single embedding for video frames to allow fast similarity-based retrieval without query-specific training. Methods such as TASTI (Kang et al., 2021) cluster frames in a learned embedding space, allowing queries to be answered using indexed embeddings rather than full model inference. While these approaches eliminate the per-query model burden, they may struggle with open-ended queries (i.e., broad or natural-language requests) if novel or uncommon concepts are poorly represented in precomputed embeddings.
Vision-Language Pretrained (VLP) embeddings. More recent approaches (Romero et al., 2023; Moll et al., 2023) provide a more general solution by leveraging VLP models like CLIP (Radford et al., 2021), which map image and text queries into a shared semantic space. These models support open-ended, natural-language-based video queries without requiring predefined object classes or embedding indexes.
However, recent works (Moll et al., 2023; Wu et al., 2024) illustrate that relying solely on models has limitations in practical video retrieval scenarios. While VLP models exhibit strong generalization capabilities, VLP models often struggle with uncommon or domain-specific queries where training data is sparse. To address these concerns, Déjà Vu leverages VideoLMs, advanced models built on top of VLP architectures that specialize in understanding video context.
2.2. Video-Language Models (VideoLMs)
VideoLMs extend VLP architectures to support complex tasks including video retrieval, video question answering, and video question grounding. For instance, video retrieval involves matching a textual description (e.g., ”Show me a clip of scenery of underwater”) to the most relevant video in a large corpus. Video question answering asks natural-language questions about the visual content, such as “How many sea turtles appear?”, requiring the model to understand both objects and context.
Computational challenges. Despite their impressive capabilities, the high computational complexity of VideoLMs remains a critical hurdle for potential applications. A major contributor to VideoLM overhead is the embedding generation process using large-scale VLP models (Yuan et al., 2021; Singh et al., 2022; Radford et al., 2021; Kim et al., 2021; Li et al., 2022b, c; Jia et al., 2021; Li et al., 2021; Sun et al., 2022; Shen et al., 2022), which commonly rely on Vision Transformer (ViT) (Dosovitskiy et al., 2021) architectures. ViTs often comprise hundreds of millions of parameters and require billions of FLOPs per inference which need to run on numerous video frames.
Figure 2. FLOPs breakdown across three VideoLM tasks: video retrieval (CLIP4Clip), video question answering (FrozenBiLM), and video question grounding (TempCLIP).
Figure 2 illustrates the computational breakdown for three VideoLMs. Notably, embedding generation via ViT dominates the overall FLOPs, underscoring the need for efficient ViT acceleration.
2.3. Vision Transformer (ViT) Acceleration
Architecture of ViT.
Figure 3. ViT model architecture.
Figure 3 illustrates ViT architecture, which adapts the transformer architecture from natural language processing to the vision domain. In ViTs, an input image or video frame is partitioned into fixed-size patches, which are linearly embedded to form a sequence of tokens. A class token (CLS) is appended to this sequence to aggregate global information. The tokens pass through Transformer encoder layers, performing query-key-value projections, multi-head self-attention, and feed-forward network.
Existing acceleration methods. To address the computational demands of ViTs, various methods exploit redundancies in the data. Acceleration techniques focusing on inter-token redundancies reduce computation by pruning less important tokens or merging similar ones within a single image (Rao et al., 2021; Fayyaz et al., 2022; Liang et al., 2022; Kong et al., 2022; Cao et al., 2023; Li et al., 2022a; Wang et al., 2021b; Dong et al., 2023; Qin et al., 2023; Bolya et al., 2023; Marin et al., 2021; Chen et al., 2023). While these pruning and merging methods effectively skip unnecessary computations within a single image or frame, they ignore the redundancy across frames. In large-scale video analytics, small frame-to-frame changes can be exploited to a much greater extent. Hence, single-image token pruning might still repeat most computations for each of these frames, making it less effective for long videos in which large swaths of patches remain nearly identical over short time intervals.
More recently, other methods have attempted to leverage inter-frame computation reuse in video data (Dutson et al., 2023; Song et al., 2024; Choi et al., 2024), offering greater computational savings. Among them, CMC (Song et al., 2024) and Eventful Transformer (Dutson et al., 2023) are most relevant to Déjà Vu, as they explicitly reuse partial computations across frames. In contrast, vid-TLDR (Choi et al., 2024) merges tokens temporally, reducing redundancy through token aggregation rather than direct computation reuse. We provide detailed comparisons against CMC and Eventful Transformer in Section 7.1.
2.4. Limitations of Video ViT Acceleration
While promising, video-targeted acceleration methods exhibit limitations that hinder their practical utility.
Challenges in balancing reuse and accuracy. Existing methods rely on manually designed strategies for computation reuse, which can make it substantially difficult to locate the optimal balance between accuracy and computational savings. For instance, Eventful Transformer (Dutson et al., 2023) requires fixing the number of tokens that reuse computation at each layer. Due to the large search space created by the tens of encoder layers and the variability in video content, it is challenging to predict the consequences of increasing reuse in certain layers. As a result, these methods may lead to a suboptimal tradeoff between computation reuse and accuracy.
Challenges in realizing FLOP savings as speedups. While reducing computational complexity (FLOPs) is crucial, achieving actual speedups requires addressing runtime factors. Mixing computed and reused tokens leads to sparse computations, causing inefficiencies on GPUs optimized for dense workloads. Our empirical analysis shows that existing video-targeted ViT acceleration techniques (Song et al., 2024; Dutson et al., 2023) deliver limited speedups due to overheads in memory usage, data movement, and hardware utilization (see Section 7). Furthermore, some prior acceleration works (Song et al., 2024) rely on specialized hardware accelerators, which are not readily available in standard commodity systems, limiting the accessibility.
These limitations motivate us to design a customized ViT model, dubbed ReuseViT, which can automatically identify computation reuse opportunities in video data, while carefully balancing the accuracy-reuse tradeoff. Additionally, we introduce memory-compute joint compaction techniques to effectively convert ReuseViT’s FLOP savings into tangible performance gains.
3. Model Architecture of ReuseViT
To address computational challenges in ViT-based VLMs, we propose ReuseViT, a model that automatically identifies safe (accuracy-preserving) computation reuse opportunities to accelerate inference. ReuseViT is designed to maximally reduce redundant computations while maintaining high accuracy by leveraging inter-frame similarities in video data.
3.1. Frame Selection for Computation Reuse
Determining which frames to reference is critical to maximizing computation reuse opportunities. In video data, especially at low frame rates common in VideoLM tasks, frame contents can drift significantly over time. Hence, ReuseViT reorders frames to allow referencing both the past and future frames, potentially boosting the likelihood of finding similar content.
Sequential frame computation.
Figure 4. (a) A basic approach where each frame references preceding frames sequentially, (b) Our proposed reordering where frames reference both past and future frames.
Figure 4(a) illustrates a basic strategy where each new frame references one or more preceding frames. Because nearer frames generally exhibit higher temporal similarity, referencing them provides reuse benefits. However, including frames further back tends to yield diminishing returns due to overlapping patches that offer little additional information.
[…]
3.2. Layer Selection for Reuse
Deciding which layers within the ViT architecture are suitable for computation reuse is crucial for maximizing efficiency gains without compromising performance.
FLOPs breakdown of ViT layers.
Figure 5. FLOPs breakdown of core computations within a single encoder layer of vision transformers at different scale.
Figure 6. Overview of ReuseViT model operations.
Figure 5 presents the FLOPs breakdown analysis results for a transformer encoder layer of three different ViT variants. The results suggest that the query-key-value (QKV) projection and the feed-forward network (FFN) are the primary consumers of computational resources in ViTs. Unlike large language models (LLMs), where the self-attention layer is the main bottleneck due to longer sequence lengths, ViTs process shorter sequences (around 256 patches). Thus, in ViTs, the self-attention layer contributes much less to the overall computational cost relative to the QKV projection and FFN layers.
Computation patterns of layers. Besides computational cost, we consider computation patterns of each layer. The QKV projection and FFN operations are applied independently to each token, without inter-token dependencies, suiting them for computation reuse. In contrast, the self-attention operation involves interactions among all tokens, making reuse more challenging and potentially affecting model output. Given these considerations, we focus on reusing computations in the QKV projection and FFN layers, targeting the most computationally intensive and token-independent operations.
3.3. Criteria for Computations Reuse
An essential aspect of ReuseViT is determining when to reuse computations for specific tokens. We introduce two key components: (1) a Decision Layer that decides whether to reuse computations, and (2) a Restoration Layer that calibrates reused computations to align with the current frame’s context. Figure 6 depicts the overall architecture of ReuseViT.
Decision layer. The decision layer identifies tokens suitable for computation reuse by assessing multiple informative cues for each token. By integrating various inputs, the layer makes nuanced decisions without relying on hand-crafted rules. The decision layer takes as input a concatenation of the following features:
  • Similarity Measure ($s$): We compute the cosine similarity between the current token and corresponding tokens in the reference frames. A higher similarity suggests that the token content has remained unchanged, making it a candidate for reuse:
    (1) $s_{i}=\max\left(\cos\left(T^{\text{cur}}_{i},T^{\text{past}}_{i}\right),\ \cos\left(T^{\text{cur}}_{i},T^{\text{future}}_{i}\right)\right)$
    Here, $T^{\text{cur}}_{i}\in\mathbb{R}^{D}$ represents the $i$-th token of the current frame, while $T^{\text{past}}_{i}$ and $T^{\text{future}}_{i}$ are the corresponding tokens from the previous and future reference frames. This strategy follows prior work on token merging using similarity metrics (Bolya et al., 2023; Dong et al., 2023; Chen et al., 2023).
  • Token Importance ($t$): We use the attention weights from the class token to estimate each token’s importance. Tokens with higher attention are more critical and may require fresh computation. This method is consistent with existing approaches that prune tokens based on class-token attention (Fayyaz et al., 2022; Liang et al., 2022; Xu et al., 2022b; Chen et al., 2023).
  • Reference Type ($r$): Reference frame type (e.g., I-frame, P-frame, Bdist1, Bdist2) offers insight into the reference’s temporal proximity, aiding assessment of reused computation reliability
  • Codec Metadata ($c$): Metadata from video codecs offers block-wise hints about spatio-temporal redundancies, providing insights into areas where video content undergoes motion or structural changes. These metadata signals, such as motion vectors and residuals, have been leveraged in prior research (Hwang et al., 2022; Yuan et al., 2023; Zhu et al., 2024) to guide computational optimizations.
By combining these inputs, the decision layer, implemented as a simple two-layer MLP, makes informed decisions.
(2) $v_{i}$ $=\text{concat}(s_{i},t_{i},r_{i},c_{i}),\quad i=1,\ldots,N$
(3) $d_{i}$ $=\text{MLP}_{\text{decision}}(v_{i})$
(4) $\mathcal{M}_{i}$ =$=\begin{cases}1,&\text{if }d_{i}>0\\ 0,&\text{otherwise}\end{cases},\quad i=1,\ldots,N$
where $\mathcal{M}_{i}\in{0,1}$ indicates whether to reuse ($1$) or recompute (00) the computations for token $i$. By combining similarity scores with token importance and codec metadata, the decision layer captures both visual and semantic cues, striking a balance between accuracy and reuse. This data-driven approach enables the model to adapt effectively to diverse video scenarios without manual tuning of thresholds or heuristics.
Token filtration. Based on the reuse map $\mathcal{M}$, tokens are partitioned into recompute tokens $C$ and reuse tokens $R$.
(5) $C$ $={T^{\text{cur}}_{i}\mid\mathcal{M}_{i}=0},\quad\text{(re\text@underline{c}ompute tokens)}$)
(6) $R$ $={T^{\text{cur}}_{i}\mid\mathcal{M}_{i}=1},\quad\text{(\text@underline{r}euse tokens)}$)
Recompute tokens $C$ undergo the standard feed forward network and query-key-value projection.
(7) $\tilde{C}^{\text{cur}}=\text{QKV}\left(\text{FFN}\left(C^{\text{cur}}\right)\right)$
Restoration layer. To adjust for discrepancies between reused computations and the current frame, we introduce a restoration layer that calibrates the reused tokens. For each reuse token $i$, we compute the difference between the tokens.
(8) $\Delta R_{i}=R^{\text{cur}}_{i}-R^{\text{ref}}_{i}$
This difference captures changes between the current and reference tokens. The restoration layer processes $\Delta R_{i}$ using a small two-layer MLP with a hidden size of 128, which is significantly smaller than the hidden size of 1,024 used in the FFN layers. This design choice reduces computational overhead while efficiently obtaining a calibration value for the change in the input.
(9) $\hat{R}^{\text{cur}}_{i}=\tilde{R}^{\text{ref}}_{i}+\text{MLP}_{\text{restoration}}(\Delta R_{i})$
where $\tilde{R}^{\text{ref}}_{i}$ is the reused computation from the reference frame, and $\hat{R}^{\text{cur}}_{i}$ is the calibrated token for the current frame. This calibration effectively improves accuracy with minimal computational overhead (4% compared to standard QKV and FFN computation).
Token Reconstruction We reconstruct the full set of tokens $\hat{T}^{\text{cur}}$ by combining the recomputed tokens $\tilde{C}^{\text{cur}}$ and the reused tokens $\hat{R}^{\text{cur}}$ according to the reuse map $\mathcal{M}$.
(10) $\hat{T}^{\text{cur}}_{i}=\begin{cases}\tilde{C}^{\text{cur}}_{j},&\text{if }\mathcal{M}_{i}=0\\ \hat{R}^{\text{cur}}_{k},&\text{if }\mathcal{M}_{i}=1\end{cases}$
where $i$ indexes the original token sequence, and $j$ and $k$ index the recomputed and calibrated reused tokens, respectively. This process maintains the original order of tokens, preserving positional relationships within the model.
The low-complexity design of the decision and restoration modules (e.g., small MLPs) ensures modest overhead relative to full inference, yielding significant net savings in large-scale deployments despite the small cost of restoring reused tokens.
4. Training ReuseViT
ReuseViT’s efficiency requires training its decision and restoration layers for a given ViT model and dataset, keeping the pre-trained ViT model frozen. This section outlines these training techniques, focusing on handling discrete reuse decisions and modeling error accumulation through grouped frame training.
4.1. Handling Discrete Reuse Decisions
A key training challenge for ReuseViT is handling discrete reuse decisions, which obstruct gradient flow through the gating mechanism during backpropagation. Specifically, the binary decisions to reuse or recompute computations for tokens hinder gradient-based optimization essential for training.
To enable gradient flow through the gating mechanism, we approximate the hard binary decisions with continuous values during training. This soft gating mechanism allows the decision layer to be trained end-to-end using backpropagation without the need for specialized techniques. By replacing the hard decisions in Equation 4 with continuous approximations, we ensure that gradients can propagate through the layer, facilitating effective learning.
(11) $\mathcal{M}_{\text{soft}}$ $=\text{GumbelSoftmax}\left(\text{MLP}_{\text{decision}}(v)\right)\approx[0,1]^{N}$
(12) $\hat{T}_{\text{soft}}^{\text{cur}}$ $=\mathcal{M}_{\text{soft}}\odot\hat{T}^{\text{cur}}+\left(1-\mathcal{M}_{\text{soft}}\right)\odot\tilde{T}^{\text{cur}}$
Here, $\odot$ denotes element-wise multiplication, $\mathcal{M}_{\text{soft}}$ is the soft mask for all $N$ tokens, $\hat{T}^{\text{ref}}$ represents the reused computations from reference frames, and $\tilde{T}^{\text{cur}}$ represents the recomputed tokens of the current frame. We found that gradually lowering the Gumbel-Softmax temperature over the course of training helps the model transition smoothly from a fully soft gating regime toward more selective token reuse. This annealing avoids sudden spikes in the reuse mask while preserving stable gradient flow.
At inference time, we apply the hard decisions to realize the computational savings achieved by ReuseViT.
4.2. Loss Function for Accuracy and Efficiency
To train ReuseViT effectively, we formulate a loss function balancing model accuracy and computational efficiency, aiming for performance gains without degradation in prediction quality.
Training objective. Our objective is to ensure that the approximated features produced by ReuseViT remain close to the original features generated by the unmodified ViT model. By preserving the original embeddings, we maintain the model’s predictive performance across different tasks without the need for task-specific fine-tuning. This self-supervised approach allows the model to generalize effectively in practical systems.
Similarity loss. We use the cosine similarity metric to encourage the approximated features to remain close to the original features of the VLP model.
(13) $\mathcal{L}_{\text{sim}}$ $=1-\cos\left(Z^{\text{cur}},\hat{Z}^{\text{cur}}\right)$
Here, $Z^{\text{cur}}$ is the original final feature vector, and $\hat{Z}^{\text{cur}}$ is the corresponding feature vector after computation reuse and calibration. However, using only the similarity loss may discourage computation reuse, as the model can minimize loss by recomputing all tokens, negating efficiency gains.
Reuse loss. To incentivize reuse, we introduce a reuse loss.
(14) $\mathcal{L}_{\text{reuse}}$ $=\frac{1}{LN}\sum_{l=1}^{L}\sum_{i=1}^{N}\mathcal{M}_{l,i}$
This loss term represents the average reuse rate across all tokens and layers. By maximizing $\mathcal{L}_{\text{reuse}}$, we encourage the model to reuse more computations, promoting computational efficiency.
Combined loss function. The overall loss function combines the similarity loss and the reuse loss:
(15) $\mathcal{L}$ $=\mathcal{L}_{\text{sim}}+\alpha\cdot\max\left(0,R_{\text{target}}-\mathcal{L}_{\text{reuse}}\right)$
Here, $\alpha$ is a weighting hyperparameter that balances the trade-off between accuracy (controlled by $\mathcal{L}_{\text{sim}}$) and computational efficiency (encouraged by $\mathcal{L}_{\text{reuse}}$), and $R_{\text{target}}$ is the target reuse rate. The $\max$ ensures that the penalty is applied only when the reuse rate falls below the target, thus incentivizing the model to meet $R_{\text{target}}$. The combined loss function effectively prevents the model from trivially minimizing error through complete recomputation.
4.3. Grouped Frame Training for Robustness
Errors accumulate over time when computations are reused across multiple frames. To ensure robustness, we model error accumulation during training by adopting a grouped frame training strategy.
Modeling error accumulation. During training, we adjust our loss functions to account for grouped frames. Specifically, we compute the losses over groups of frames by averaging the similarity loss $\mathcal{L}_{\text{sim}}$ and the reuse loss $\mathcal{L}_{\text{reuse}}$ across all frames in the group. By considering the average losses over the group, the model learns to handle error accumulation effectively over sequences of frames, rather than optimizing for individual frames in isolation. This approach ensures that the model’s performance and reuse behavior are optimized not just for single frames, but for sequences where errors may propagate due to computation reuse.
Efficient grouping strategy. To balance modeling accuracy and training efficiency, we design training sequences that model error accumulation over a manageable number of frames.
Figure 7. Illustration of our frame-grouping strategy during training. We treat each “I” or “P” frame as a segment boundary and form groups allowing the model to learn error accumulation.
We use I-frames and P-frames as segment boundaries, dividing frames into segments (e.g., Segment 1 and Segment 2 in Figure 7). Frames within a segment can influence each other through computation reuse, while intermediate frames across segments remain independent, allowing for efficient modeling of longer sequences.
By constructing sequences like I-P-P, corresponding to frames 1, 5, and 9, we simulate error propagation over longer intervals without including intermediate frames (2–4 and 6–8). Within the last segment, we also include frames with remaining Bdist2 and Bdist1 reference types from the previous segment, so that ReuseViT is able to learn all reference types.
Empirical results indicate that grouping six frames in the pattern 1-5-9-13-11-12 strikes an effective balance between accuracy and training efficiency, allowing the model to account for error accumulation over significant temporal spans while managing computational costs. During training, we also experimented with varying group sizes and found that 6 frame groupings gave us the best empirical tradeoff between training time and error-accumulation robustness. Larger groupings occasionally made optimization less stable, while smaller ones did not fully capture longer-range drifts.
5. Compaction Techniques in Déjà Vu
In this section, we present the compaction methods used to translate the FLOPs reduction achieved by ReuseViT into tangible performance gains. These methods address practical challenges in implementing ReuseViT efficiently on GPU hardware, where parallelism is crucial to achieving high throughput.
5.1. Layer-Wise Scheduling
A key enabler of our compaction methods is layer-wise scheduling. While the target problem differs, our scheduling scheme is inspired by FlexGen (Sheng et al., 2023), a prior work that aims to deploy a large language model (LLM) serving on a single GPU. FlexGen proposes grouping multiple batches for LLM inference and iterates the batches in a round-robin to execute their computations layer-by-layer. That way, FlexGen can amortize the I/O costs across multiple batches, required for model parameters and KV cache.
On the other hand, Déjà Vu has different objectives since it aims to accelerate ViT encoders, which have smaller model sizes than LLMs and do not need KV cache. Déjà Vu employs the layer-by-layer computation scheme to mitigate the memory bloating and GPU unfriendliness problems caused by ReuseViT’s computation reuse. Instead of interleaving multiple batches as in FlexGen, Déjà Vu batches multiple frames together and processes the same layer across these frames in a layer-by-layer manner within the same segment. Building upon this layer-wise scheduling, Déjà Vu employs Cached Memory Compaction and Sparse Computation Compaction to more efficiently utilize GPU memory and compute resources.
5.2. Cached Memory Compaction
A major overhead in computation reuse arises from storing cached activations, or intermediate outputs computed for reference frames, which subsequent frames rely on.
Layer-wise memory compaction. By exploiting our layer-wise scheduling and the frame-referencing structure of our model (with P-frames acting as boundaries), we compact memory usage at each layer within a segment. Frames other than P-frames do not affect future segments, so their intermediate activations can be discarded once their current segment is processed.
Figure 8. Comparison of (a) frame-wise scheduling and (b) layer-wise scheduling for caching intermediate activations.
Figure 8 illustrates the key difference between the conventional frame-wise approach. As illustrated in Figure 8(a), in the conventional frame-wise approach, each frame runs through all layers sequentially. Intermediate outputs must remain in memory until all computations for that frame are complete. In contrast, in our layer-wise scheduling (Figure 8(b)), the same layer is applied to all frames in the batch, and once a layer is complete for each frame, any unneeded activations are immediately freed. This staggered approach substantially lowers memory overhead, enables larger batch sizes with high GPU utilization, and facilitates deployment of large-scale VideoLMs under limited memory budgets.
5.3. Sparse Computation Compaction
ReuseViT’s dynamic and often sparse computation patterns pose difficulties for GPUs, which are optimized for dense, regular workloads. To address this, we use stream compaction (Billeter et al., 2009), gathering active (i.e., non-reused) tokens into contiguous memory regions so sparse computations can be converted into dense forms better suited to GPU acceleration.
Figure 9. Sparse Computation Compaction.
Layer-wise stream compaction. We implement GPU kernels for stream compaction across different frames within the same segment, as shown in Figure 9. By accumulating active tokens from multiple frames before moving on to the next feed forward network, we form well-shaped matrices amenable to GPU acceleration. Once the batched computation is complete, we scatter the results back into their original positions in each frame’s token sequence, preserving correctness for the next layer. This gather-scatter strategy ensures that even when some frames have only a few active tokens, they can still benefit from dense GPU operations.
To minimize overhead further, we prioritize CUDA-based implementations for the gather-compute-scatter procedure, avoiding frequent CPU-GPU synchronization. Such an approach not only speeds up the compaction process itself but also reduces latency spikes that might occur from excessive kernel launches or data transfers between CPU and GPU.
6. Workflow of Déjà Vu
6.1. Overview of Query Processing
On a query, Déjà Vu returns cached embeddings if available; otherwise it generates them with ReuseViT and stores the result.
Embedding storage overhead. Storing frame-level VLP embeddings introduces minimal storage overhead. For example, extracting embeddings at 2 FPS from a Full-HD H.264 video ($\sim$625KB/s) results in approximately 4KB/s of data (assuming 1024-dimensional FP16 embeddings at $\sim$2KB each). This constitutes merely 0.64% of the compressed video size, with further footprint reductions achievable via advanced floating-point compression techniques (Liakos et al., 2022; Pelkonen et al., 2015; Li et al., 2023a).
6.2. Offline Preparation
Before handling real queries, Déjà Vu trains the decision and restoration modules for a given ViT backbone. During training, we use the self-supervised objective, which combines a similarity loss (to match the original ViT outputs) and a reuse-based loss (to encourage high reuse rates). The primary user-adjustable hyperparameter is the target reuse rate ($R_{\text{target}}$), defined in Equation 15. Training with different values of $R_{\text{target}}$ allows users to navigate various points on the accuracy-speed tradeoff curve. Alternatively, one can specify a target cosine similarity, whereby the model learns to maximize the reuse rate while conforming to the accuracy threshold defined by that similarity value.
We also employ grouped-frame training to model how errors might accumulate. This improves robustness, as the model learns to balance computational savings with accuracy preservation across different video types. During training, Only the two lightweight modules are trained, while the pre-trained ViT backbone remains frozen and convergence typically occurs within an hour. Such design significantly simplifies deployment by removing the need to store, transfer, or manage multiple versions of large model weights.
6.3. Online Inference
Once deployed, Déjà Vu employs layerwise scheduling and memory compaction to translate the FLOP savings from ReuseViT into practical GPU throughput. While ReuseViT reduces the token-level computational load, effective scheduling and compaction routines ensure these savings materialize in actual GPU execution time which is a key requirement for large-scale query systems.
To maximize GPU utilization, the system processes multiple videos in a single batch. If only one long video is available, it is split into multiple, uniformly sized segments. For each segment, four consecutive frames are collected via a small queue and fed into ReuseViT, enabling each inference call to process four times the usual batch size in frames. Batching multiple segments is crucial for efficiently utilizing GPU resources, especially when high reuse rates substantially reduce computation per video.
Frame reordering. Frame reordering is handled within the model’s forward pass, which reconstructs results in the correct sequence. Consequently, the outer framework does not need to manage any frame shuffling or realignment. Although this staggered approach works well for offline analytics, where latency is less critical, it can also be adapted to streaming scenarios. In a streaming setting, Déjà Vu buffers four frames before computation, introducing a queuing latency of about two seconds at 2 FPS. For real-time applications with strict latency requirements, frame reordering can be disabled to minimize latency, albeit at the cost of lower reuse rates.
Breaking error propagation. Although grouping frames largely confines errors, prolonged reuse can still amplify minor inaccuracies. To counter this, Déjà Vu periodically inserts a fully recomputed I frame. For instance, if the system resets every twentieth frame, it prevents indefinite error buildup without incurring excessive recomputation costs. This method keeps the overhead below 5% while maintaining accuracy over long sequences.
By blending standard ViT-based embedding generation with lightweight, learned modules for inter-frame reuse, Déjà Vu delivers on-demand embeddings for a variety of queries. Its approach avoids repeated backbone retraining, recycles partial results across neighboring frames, and capitalizes on compaction techniques to maximize GPU efficiency. These design choices not only make Déjà Vu robust to queries demanding high accuracy but also ensure practical deployment for both offline and streaming scenarios. As a result, Déjà Vu opens the door to scalable, open-ended video-language applications without the prohibitive compute costs usually associated with large-scale video analytics.
7. Evaluation
Figure 10. Tradeoff space of baselines and Déjà Vu. Y-axis for FLOPs reduction and throughput are normalized to original model with no reuse.
7.1. Methodology
End tasks. To evaluate the effectiveness of Déjà Vu, we select three distinct Video-Language Models (VideoLMs), each targetting a different task. All these models use embeddings extracted from CLIP (Radford et al., 2021) (ViT) as their visual backbone. We focus on:
  • Video retrieval. We use CLIP4Clip (Luo et al., 2022) for video retrieval and evaluate on MSR-VTT (Xu et al., 2016) dataset, which consists of 10,000 video clips annotated with textual descriptions, typically 10 to 30 seconds long. We report top-5 recall as the accuracy metric.
  • Video question answering. We use FrozenBiLM (Yang et al., 2022a) for video question answering (or video QA), a task where the goal is to answer queries based on video content. We evaluate on How2QA (Li et al., 2020a), which standardizes input to 60 second clips drawn from longer instructional videos with 44,007 question-answer pairs, and report multiple-choice accuracy as the metric.
  • Video Question Grounding. Video question grounding (or video grounded QA) requires the model to predict the relevant video segment along its answer. We use TempCLIP (Xiao et al., 2024) and evaluate it on NExT-GQA (Xiao et al., 2024), with longer, untrimmed videos averaging around 45 seconds, with short answer segments requiring fine-grained temporal grounding. We report GQA@Accuracy, the percentage of correct answers where the predicted video segment overlaps the ground-truth segment.
To ensure a fair comparison, all baselines sample frames at 2 frames per second, which aligns with common practice in recent video-language models (Piergiovanni et al., 2023; Arnab et al., 2021; Bulat et al., 2021; Zhang et al., 2023; Wang et al., 2024a; Luo et al., 2022; Yang et al., 2022a; Xiao et al., 2024). Throughout all tasks, we use the standard train, val, and test splits provided by each dataset and adhere to the recommended evaluation protocols.
Baselines. As no prior methods directly accelerate VideoLMs, our baselines include temporal ViT acceleration approaches (conceptually closest to Déjà Vu’s strategy) and a representative image-based method for contrast. We evaluate three acceleration methods:
  • CMC. CMC (Song et al., 2024) uses a fixed threshold value on the mean squared error (MSE) to identify unchanged tokens suitable for reuse across frames. It leverages similarity computations provided by specialized video codec hardware and was originally evaluated on action recognition tasks.
  • Eventful Transformer. Eventful Transformer (Dutson et al., 2023) applies a static, layer-wise policy that fixes the number of tokens to recompute, regardless of changes in video content. While simple and efficient, this rigid strategy might not adapt effectively to varying temporal content across frames, potentially causing error accumulation in dynamic scenes.
  • DiffRate. DiffRate (Chen et al., 2023) is a representative baseline for images, combining token pruning and merging within an image. It automatically determines the tradeoff between accuracy and efficiency through backpropagation. Though originally evaluated on image classification, we adapted the policy for VLP models.
CMC and Eventful Transformer propose kernels to reuse computations in the attention layer. In our experiments, these kernels did not yield speedups because the attention layers have relatively small computational costs. Therefore, for both methods, we focus on their core reuse strategies without attention-layer optimizations.
Implementation details. We run all experiments on a server with two 16-core Intel Xeon Gold 6226R CPUs (2.9 GHz), 192 GB of RAM, and an NVIDIA RTX 3090 GPU (24 GB of GDDR6 memory). The software environment includes Ubuntu 24.04 as the operating system, CUDA 12.1, cuDNN 8.9.0, and PyTorch 2.1.0.
7.2. Tradeoff between FLOPs and Accuracy
Figure 10 (a)–(c) shows tradeoffs between FLOPs reduction for embedding generation (y-axis) and accuracy (x-axis) for three tasks: video retrieval, video QA, and video question grounding. The x-axis shows the accuracy metric, and the star-shaped marker near the bottom right of each graph indicates the unapproximated model’s accuracy. The y-axis shows the achieved FLOPs reduction for embedding generation. A point higher on the y-axis indicates greater FLOPs reduction at the same accuracy level, while a point further to the right indicates better accuracy. Thus, methods near the top-right corner represent more favorable efficiency-accuracy tradeoffs. Next to each marker, we show the computation reuse rate, except for DiffRate, which applies token pruning and merging instead.
Notably, techniques that exploit temporal redundancy (CMC, Eventful Transformer, and Déjà Vu) consistently achieve higher FLOPs reduction than the image-focused DiffRate, whose pruning/merging targets only spatial redundancy within individual frames. Among these temporal-based baselines, Déjà Vu attains the best tradeoff, typically matching or exceeding each method’s accuracy at a lower computational cost.
7.3. Tradeoff between Throughput and Accuracy
Figures 10 (d)-(f) show throughput improvement for the embedding generation on the y-axis. While Déjà Vu and DiffRate are measured on a real GPU system, CMC and Eventful Transformer lack GPU-optimized implementations, so we interpolate their throughput assuming our compaction techniques applied. Here, we normalize throughput to that of the original model, and only configurations that yield an actual improvement are shown, and each marker can be matched to those in the FLOPs plots via the reuse rates.
Consistent with previous FLOPs-based results, Déjà Vu achieves the highest throughput-accuracy tradeoff. For instance, Déjà Vu reaches speedups of up to 1.81$\times$ for video retrieval, 2.64$\times$ for video question answering, and 2.54$\times$ for video question grounding. By contrast, Eventful Transformer, the next-best baseline, achieves speedups of 1.32$\times$, 2.08$\times$, and 2.20$\times$, respectively.
While baselines exploiting inter-frame redundancies remain competitive for VideoLM tasks, Déjà Vu’s tailored optimizations achieve superior throughput-accuracy tradeoffs. Even incremental improvements are crucial, for instance, a 1–2% accuracy gain in leading video QA benchmarks typically requires several months of intensive research (Yang et al., 2022a, 2021), and modest speedups can significantly reduce costs and query latency in large-scale deployments.
Interestingly, DiffRate configurations become more competitive for throughput, as the high complexity of computation reuse can hinder its realization. However, Déjà Vu’s careful optimizations still ensure higher overall speedups.
7.4. Breakdown Analysis
At the same reuse rate, Déjà Vu has slightly lower performance than CMC or Eventful Transformer. For instance, in Figure 10 (e) when reuse is at 61%, Déjà Vu shows a marginally lower throughput.
Figure 11. Breakdown of FLOPs on video retrieval.
Figure 11 (a) shows a FLOPs breakdown of these methods at an identical reuse rate on the video retrieval task. In this scenario, the reuse rate is intentionally fixed across all methods to isolate and highlight algorithmic overhead. The main additional overhead in ReuseViT arises from the decision layers and restoration layers, which add about 4% to normalized FLOPs.
However, Figure 11 (b) shows that at the same accuracy on video question answering, Déjà Vu achieves a significantly higher reuse rate. Thus, at equivalent accuracy levels, Déjà Vu effectively compensates for the additional overhead by adaptively calibrating reused computations, enabling it to reuse more tokens without sacrificing accuracy. Hence, while Déjà Vu incurs some overhead at the same reuse rate, its adaptive mechanisms enable higher reuse where accuracy must be maintained, which ultimately yields stronger end-to-end speedups when matching target accuracy levels.
7.5. Memory Overhead
Figure 12. Comparison of peak memory usage with and without cached memory compaction. The dotted line shows memory capacity of RTX 3090 GPU.
Figure 12(a) compares the peak GPU memory usage of ReuseViT with and without cached memory compaction, across different batch sizes. The blue bars show memory usage when intermediate computations are cached without compaction, while the green bars include compaction. Without compaction, the model processes only 67 batches before running out of memory, but compaction raises this limit to 183, improving GPU utilization and throughput. By keeping memory usage within GPU limits, practitioners can deploy larger batch sizes or process multiple video streams concurrently, which in turn boosts overall system throughput. This is particularly valuable in server-based or cloud environments where memory is a critical and often expensive resource.
Figure 12(b) illustrates how the cache size changes during inference. Without compaction, the cache grows linearly until all layers are processed. With compaction, cached memory for processed frames is freed after certain layers, creating a sawtooth pattern and significantly reducing peak memory usage.
7.6. Ablation Study for Inference Speedup
Figure 13. Ablation analysis compaction techniques.
We conduct an ablation study of inference speedup with a ReuseViT configuration achieving a 61% reuse rate on the video question answering task to show how each technique contributes to Déjà Vu ’s performance. Figure 13 highlights the throughput gains from each step. The first bar indicates reusing computation with hard gating alone yields a 1.25$\times$ speedup. Layer-wise sparse computation compaction then increases the speedup to 1.45$\times$ by shaping operands to better utilize the GPU kernel. Finally, layer-wise memory compaction enables larger batches, improving parallelization and raising the total speedup to 1.62$\times$. These results confirm Déjà Vu’s optimizations translate FLOPs savings into real throughput gains. Each component offers complementary benefits, gating reduces compute, sparse-compaction improves GPU workload efficiency, and memory compaction allows larger batches, unlocking significantly more performance together than individually.
7.7. Adaptability to Video Content
We compare the adaptability of Eventful Transformer and Déjà Vu on a video segments from How2QA, with both methods reaching 84.5% accuracy.
Figure 14. Fluctuation of computation reuse rate and cosine similarity over time.
Figure 14 depicts the fluctuations in computation reuse rates (top) and cosine similarities (bottom) over time.
Initially, both methods fully compute the first frame, then adopt their reuse strategies. Eventful Transformer maintains a fixed reuse rate, leading to a more pronounced decline in cosine similarity. In contrast, Déjà Vu adapts when error begins to accumulate, lowering the reuse rate to protect accuracy and raising it when frames are more similar. Moreover, Déjà Vu learns to adjust reuse behavior based on reference frame types (I- or P-frames), further mitigating error accumulation. By dynamically balancing reuse and recomputation, Déjà Vu preserves higher feature similarity over time and achieves a superior tradeoff between computational efficiency and model accuracy compared to methods with fixed reuse strategies.
7.8. Ablation Study for Design Choices
To evaluate the contributions of ReuseViT’s core components, we perform an ablation study on the NeXT-GQA dataset.
Figure 15. Ablation study of ReuseViT’s design choices on the NeXT-GQA dataset. The full ReuseViT configuration is denoted by a triangular marker.
Figure 15 assesses visual embedding quality, measured by cosine similarity to the original embedding without reuse, and throughput improvement normalized against the original ViT. Higher values on both axes indicate better efficiency and accuracy.
Figure 15(a) presents incremental component comparisons. Starting from a configuration with only the adaptive decision layer, we sequentially incorporate frame reordering, grouped frame training, and finally, the restoration layer. Notably, even the initial decision layer alone already provides substantial improvements over the static Eventful Transformer baseline, highlighting the significant advantage of adaptive reuse decisions. Each subsequent addition further enhances the tradeoff between efficiency and accuracy. Although the restoration layer introduces a slight computational overhead reducing peak throughput, it notably improves embedding quality at higher reuse rates.
Figure 15(b) examines various frame grouping strategies. Compared to a sequential processing baseline, all reordered groupings achieve improved tradeoffs. Performance consistently increases with segment lengths up to three segments, while extending to four segments yields slightly inferior results. Thus, a three-segment grouping strategy effectively balances computational efficiency and visual embedding quality.
8. Limitations and Future Work
While Déjà Vu largely improves computational efficiency for video-language tasks, several limitations remain, motivating future work.
Attention Layer Reuse. Déjà Vu currently targets only FFN and QKV layers. These layers dominate computational costs at standard resolutions, for instance, 257 tokens for ViT-L/14. However, as resolution and token counts rise—such as 577 tokens (CLIP ViT-L/14@336px), 785 tokens (DINO ViT-B/8) (Caron et al., 2021), and 1370 tokens (DINOv2 ViT-G/14@518px) (Oquab et al., 2024)—attention layers form an increasingly substantial fraction of total FLOPs, reaching up to 23.5%. Addressing this will necessitate exploring attention reuse strategies or leveraging sparse attention kernels optimized for GPUs, like those proposed in Eventful Transformer. Another approach involves specialized hardware accelerators, such as those introduced in CMC.
Broader task generalization. This work evaluates Déjà Vu specifically on retrieval, question-answering, and grounding tasks. Adapting our learned reuse strategies to other video tasks, such as action recognition or object detection, would involve adjustments to task-specific model architectures and setups. While this adaptation is conceptually straightforward, its practical exploration remains an open and promising direction for future research.
Adaptive periodic refresh and online adaptation. Another limitation relates to the adaptive periodic refresh strategy within Déjà Vu. This strategy could benefit from additional refinement for improved handling of long-form video inputs and enhanced robustness across various video domains. Furthermore, future research might explore online fine-tuning methods designed to dynamically adjust reuse policies during inference. Such adjustments would enhance adaptability and efficiency within evolving real-world scenarios.
9. Related Work
CNN-based pipelines for known classes. Early systems for known object categories often relied on trained CNNs or proxy models. For instance, NoScope (Kang et al., 2017) developed lightweight cascades for static-camera queries. Focus (Hsieh et al., 2018) and BlazeIt (Kang et al., VLDB) built approximate indexes that were refined by heavier models. Other approaches included Tahoma’s (Anderson et al., 2019) use of cascades with input transformations and MIRIS (Bastani et al., 2020) targeted multi-object tracking. Some systems precomputed object tracks, like OTIF (Bastani and Madden, 2022), or used detector ensembles, as seen in FiGO (Cao et al., 2022). More recent contributions include DoveDB (Xiao et al., 2023) unifying training and tracklet ingestion, Seiden (Bang et al., 2023) leveraging high-accuracy models with sampling to avoid proxies, and InQuest (Russo et al., 2023) paring cheap proxies with oracles for streaming analytics. While cost-effective for predefined categories, these systems struggle with zero-shot or open-ended queries without retraining.
Resource management and domain-specific analytics. Another line of work focuses on scheduling or adapting analytics pipelines at scale. Scanner (Poms et al., 2018), for example, exploits domain metadata for raw video decoding and dataflow. VideoStorm (Zhang et al., 2017) and Chameleon (Jiang et al., 2018) schedule model cascades for concurrent streams. ODIN (Suprem et al., 2020) monitors domain drift to retrain models for known classes under new conditions. Further techniques involve adaptive query reordering by Hydro (Kakkar et al., 2024), using geospatial metadata to skip frames as in Spatialyze (Kittivorawong et al., 2024), or integrating caching to accelerate iterative queries, demonstrated by VIVA (Kang et al., 2022).
Open-vocabulary indexing and domain-agnostic exploration. To address fixed label sets, several systems enable querying arbitrary classes without retraining. Panorama (Zhang and Kumar, 2020) learns a generic feature space. TASTI (Kang et al., 2021) reuses a learned semantic index for new labels, and Boggart (Agarwal and Netravali, 2023) provides a model-agnostic index. Other systems support interactive exploration, such as EVA (Xu et al., 2022a). VOCALExplore (Daum et al., 2023) trains detectors on-the-fly via active labeling, while SeeSaw (Moll et al., 2023) leverages CLIP embeddings for zero-shot retrieval. LVS (Lee and Park, 2024) approximates embeddings from cached ones, and SketchQL (Wu et al., 2024) uses sketch-based queries. Domain-agnostic storage solutions like VStore (Xu et al., 2019), TASM (Daum et al., 2021), and VSS (Haynes et al., 2021), along with decoding-stage acceleration in CoVA (Hwang et al., 2022) and TVM (Zhong et al., 2023), also enhance query speed.
Modified architectures and acceleration for ViTs. Following ViT’s (Dosovitskiy et al., 2021) success, many subsequent architectures aim to reduce computation using techniques such as local attention or convolutions (Liu et al., 2021; Wang et al., 2021a, 2022; Chu et al., 2021; Pan et al., 2022; Heo et al., 2021; Graham et al., 2021). For video processing, temporal mixing layers are common additions (Wang et al., 2024b; Cheng et al., 2023; Yang et al., 2022b; Ren et al., 2023; Tong et al., 2022; Wang et al., 2023; Li et al., 2023b), though no standard has emerged. Standard ViTs are accelerated via token pruning (removing less important tokens) (Rao et al., 2021; Fayyaz et al., 2022; Liang et al., 2022; Kong et al., 2022; Chen et al., 2023; Cao et al., 2023) or token merging (combining similar tokens) (Bolya et al., 2023; Marin et al., 2021; Chen et al., 2023; Cao et al., 2023; Li et al., 2024; Choi et al., 2024). Pruning techniques can be fixed-rate (Rao et al., 2021) or adaptive (Fayyaz et al., 2022), sometimes combined with merging strategies (Liang et al., 2022; Kong et al., 2022). Merging approaches, exemplified by ToMe (Bolya et al., 2023) and TCFormer (Marin et al., 2021), downsample tokens, with recent extensions to temporal attention (Li et al., 2024; Choi et al., 2024). Some methods integrate both pruning and merging (Chen et al., 2023; Cao et al., 2023) or propose specialized hardware (Li et al., 2022a; Wang et al., 2021b; Dong et al., 2023; Qin et al., 2023; Yoo et al., 2024).
Computation reuse and decision making. Reusing computations across frames offers gains beyond intra-frame redundancy. Eventful Transformer (Dutson et al., 2023) and CMC (Song et al., 2024) use manual cross-frame reuse decisions. In contrast, Déjà Vu employs a fully learnable mechanism for improved efficiency. Prior work on temporal redundancy in CNNs includes incremental activation updates (Buckler et al., 2018; Pan et al., 2021), use of motion vectors (Song et al., 2020), delta updates (Parger et al., 2022, 2023), or discarding redundant frames and regions (Taranco et al., 2023; Dai et al., 2022; Li et al., 2020b; Wu et al., 2019; Wakatsuki et al., 2021). Several studies also learn reuse decisions end-to-end. These include methods for dynamic fusion (Zhang et al., 2020b) or layer skipping (Wu et al., 2018; Wang et al., 2018; Chen et al., 2020). While these predominantly target CNNs, Déjà Vu adapts this concept to VideoLMs by integrating a learnable reuse mechanism in ViTs for significant speedups while preserving accuracy.
10. Conclusion
This paper addresses the computational challenges posed by ViT in the context of large-scale video analytics exploiting VideoLM. We introduce Déjà Vu, a novel algorithm-system co-designed solution, recognizing the untapped potential of exploiting a learning approach to enable inter-frame computation reuse. Central to Déjà Vu is the ReuseViT model, innovatively adapted for VideoLM tasks. This model is specifically designed for enabling computation reuse opportunities in ViT inferencing by judiciously determining the reuse targets and restoring the reuse-introducing errors automatically. We also build a ReuseViT-based Déjà Vu system, which effectively utilizes GPUs for high throughput ViT inferencing for VideoLM queries. Our evaluation shows that Déjà Vu achieves a substantial performance boost while marginally compromising the accuracy of VideoLM tasks. These findings suggest that this work represents a practical initial step toward actualizing the full potential of VideoLMs on video-language query systems.